OpenStudy (ttop0816):

cos870° - sin780° = _____ -sq.root 3 sq. root 3 0

OpenStudy (anonymous):

cos (870) = cos (150) = - sqrt(3)/2 sin (780) = sin (60) = sqrt(3)/2 So, -sqrt(3)/2 - sqrt (3)/2 = -2sqrt(3)/2 = - sqrt(3) Note: an angle of 870 = an angle of 150, because every 360 it is periodic.

OpenStudy (ttop0816):

thank you so much! what about this one?? sin400° = _____ -sin40° sin40° <-- would it be this? sin(-40)°

OpenStudy (ttop0816):

@Easyaspi314

OpenStudy (anonymous):

sin (400) = sin (40) Just subtract 360 from 400. Both angles are identical.

OpenStudy (anonymous):

And you can easily verify this on your calculator: sin(400) = .6427876097 sin (40) = .6427876097

OpenStudy (ttop0816):

so it would be the answer i chose right??

OpenStudy (ranga):

Yes the answer you have indicated is the correct one.